On May 1, 2018, Benz’s Sandwich Shop loaned $18,000 to Mark Henry for one year at 6 percent interest.


Required

  1. What is Benz’s interest income for 2018? (Round your answer to the nearest dollar amount.)
  1. What is Benz’s total amount of receivables at December 31, 2018? (Round your answer to the nearest dollar amount.)
  1. How will the loan and interest be reported on Benz’s 2018 statement of cash flows?
  1. What is Benz’s interest income for 2019? (Round your answers to the nearest dollar amount.)
  1. What is the total amount of cash that Benz’s will collect in 2019 from Mark Henry? (Round your answer to the nearest dollar amount.)
  1. How will the loan and interest be reported on Benz’s 2019 statement of cash flows? (Round your answer to the nearest dollar amount.)
  1. What is the total amount of interest that Benz’s earned on the loan to Mark Henry? (Round your answer to the nearest dollar amount.)
Answer #1

Answer

  • Working calculation

Transaction

Cash

Notes receivables

Interest receivables

Interest revenue

01-May-18

Amount lent

($18,000)

$18,000

31-Dec-18

8 month Interest income accrued

$720

$720

31 Dec 2018 Balances

($18,000)

$18,000

$720

$720

01-May-19

Accrued interest received

$720

($720)

01-May-19

Remaining 4 month Interest received

$360

$360

Principal amount received

$18,000

($18,000)

31 Dec 2019 Balances

$1,080

$0

$0

$360

Transaction

Cash

Notes receivables

Interest receivables

Interest revenue

01-May-18

Amount lent

-18000

18000

31-Dec-18

8 month Interest income accrued

=18000*6%*8/12

=18000*6%*8/12

31 Dec 2018 Balances

($18,000)

$18,000

$720

$720

01-May-19

Accrued interest received

720

-720

01-May-19

Remaining 4 month Interest received

=18000*6%*4/12

360

Principal amount received

18000

-18000

31 Dec 2019 Balances

$1,080

$0

$0

$360

  • Requirements

a

$720

b

$18,720

c

Statement of Cash Flows - Partial

Cash flows from Investing activity:

Cash loaned

($18,000)

d

$360

e

$19,080

[$ 18000 + 720 + 360]

f

Statement of Cash Flows - Partial

Cash flows from Operating activity:

Cash received for interest

$1,080

Cash flows from Investing activity:

Cash received of principal

$18,000

g

$1,080

[$ 18000 x 6%]